# Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC

### v2

@@ II-C Multi-Static Sensing
In the presence of the target, each receive AP captures both the reflected signals from the target and undesired signals, referred to as clutter. The clutter, being independent of the target’s presence, is treated as interference for sensing purposes. Without loss of generality, we assume that the LOS paths between transmit and receive APs are known and can be effectively canceled out. Consequently, the interference signals correspond to the reflected paths through obstacles and are henceforth denoted as target-free channels.
@@
Let $\textbf{H}_{r,k}\in\mathbb{C}^{M\times M}$ denote the target-free channel matrix between transmit AP $k$ and receive AP $r$. We use the correlated Rayleigh fading model for the NLOS channels $\textbf{H}_{r,k}$, which is modeled using the Kronecker model [38]. We define the random matrix $\textbf{W}_{{\rm ch},(r,k)}\in\mathbb{C}^{M\times M}$ with i.i.d. entries with $\mathcal{CN}(0,1)$ distribution. The matrix $\textbf{R}_{{\rm rx},(r,k)}\in\mathbb{C}^{M\times M}$ represents the spatial correlation matrix corresponding to receive AP $r$ and with respect to the direction of transmit AP $k$. Similarly, $\textbf{R}_{{\rm tx},(r,k)}\in\mathbb{C}^{M\times M}$ is the spatial correlation matrix corresponding to transmit AP $k$ and with respect to the direction of receive AP $r$. The channel $\textbf{H}_{r,k}$ is written as
$\textbf{H}_{r,k}=\textbf{R}^{\frac{1}{2}}_{{\rm rx},(r,k)}\textbf{W}_{{\rm ch},{(r,k)}}\left(\textbf{R}^{\frac{1}{2}}_{{\rm tx},(r,k)}\right)^{T},$ (13)
where the channel gain is determined by the geometric path loss and shadowing, and is included in the spatial correlation matrices. The received signal at AP $r$ in the presence of the target and for $m\in\{1,\cdots,L_{d}\}$, can be formulated as
@@
$+\sum_{k=1}^{N_{\rm tx}}\underbrace{\textbf{H}_{r,k}\textbf{x}_{k}[m]}_{\textrm{undesired signals}}+\textbf{n}_{r}[m],$ (14)
@@
Following the same notation as in [34], in (II-C), the known part of each reflected path is denoted by ${\textbf{g}_{r,k}[m]\in\mathbb{C}^{M}}$, defined as
@@
Each receive AP sends their respective signals $\textbf{y}_{r}[m]$, for $r=1,\ldots,N_{\rm rx}$, to the edge cloud, to form the collective received signal $\textbf{y}[m]=\begin{bmatrix}\textbf{y}_{1}^{T}[m]&\cdots&\textbf{y}_{N_{\rm rx}}^{T}[m]\end{bmatrix}^{T}$, which can be expressed as
@@
$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\bm{\mathfrak{h}}+\textbf{n}[m],$ (16)
where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\cdots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\cdots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\mathfrak{h}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channels[34, Sec. V].
II-D Target Detection
We employ two maximum a posteriori ratio test (MAPRT) detectors with two levels of complexity: the clutter-unaware detector [13] and the clutter-aware detector [34, Lem. 2]. The clutter-unaware detector ignores the presence of clutter for the sake of complexity reduction, whereas the clutter-aware detector accounts for unknown clutter. The corresponding test statistics are given by
@@ II-D Target Detection
II-E Transmit ISAC Precoding Vectors
where $\delta$ is the regularization parameter, and $\hat{\textbf{h}}_{j}=\begin{bmatrix}\hat{\textbf{h}}_{j,1}^{T}&\ldots&\hat{\textbf{h}}_{j,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the LMMSE channel estimate of the communication channel $\textbf{h}_{j}$.
and U is the unitary matrix with the orthogonal columns that span the column space of the matrix $\begin{bmatrix}\hat{\textbf{h}}_{1}&\ldots&\hat{\textbf{h}}_{N_{\rm ue}}\end{bmatrix}$. $\textbf{h}_{0}=\begin{bmatrix}\sqrt{\beta_{1}}\textbf{a}^{T}(\varphi_{1},\vartheta_{1})&\ldots&\sqrt{\beta_{N_{\rm tx}}}\textbf{a}^{T}(\varphi_{N_{\rm tx}},\vartheta_{N_{\rm tx}})\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the concatenated sensing channel between all the ISAC APs and the target, including corresponding path loss and the array response vectors $\textbf{a}(\varphi_{k},\vartheta_{k})$ for $k=1,\ldots,N_{\rm tx}$.
@@ III Reliability and Delay Analysis for URLLC
In this section, we derive an upper bound on the DEP and the transmission delay, both crucial aspects considered as URLLC requirements. In the finite blocklength regime, the communication data cannot be transmitted without error. From [10], ergodic data rate of UE $i$ can be approximated as
$R_{i}\!\approx\!\mathbb{E}\Bigg{\{}\!\!\left(1\!-\!\beta\right)\!\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\!\sqrt{\!\frac{(1\!-\!\beta)V_{i}}{L}}\Bigg{\}},\vspace{-0mm}$ (25)
where $\beta=\frac{L_{p}}{L}$, $\epsilon_{i}$ denotes the DEP when transmitting $b_{i}$ bits to UE $i$, $\mathsf{SINR}^{\rm(dl)}_{i}$ is the instantaneous downlink communication signal-to-interference-plus-noise ratio (SINR) for UE $i$, $V_{i}=1-\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)^{-2}$ is the channel dispersion, and $Q(\cdot)$ refers to the Gaussian Q-function. Due to the fact that $V_{i}<1$, the ergodic data rate can be lower bounded by
@@
Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [35, Thm. 6.1] and [34, Lem. 1],
@@
where $D_{i}^{\rm(th)}$ is the maximum tolerable delay by UE $i$ and $D_{i}^{\rm(ub)}\leq D_{i}^{\rm(th)}$ should be satisfied to guarantee the delay requirement. This implies that the blocklength cannot exceed $D^{\rm(th)}_{i}B(1-\epsilon_{i}^{\rm(th)})$. Thus, we can define the maximum tolerable blocklength by $L_{\rm max}$, where
$L_{\rm max}=\min\left\{D^{\rm(th)}_{i}B\left(1-\epsilon_{i}^{\rm(th)}\right)|i=1,\cdots,N_{\rm ue}\right\}.$ (34)
@@ IV Sensing Analysis
We assess the sensing performance in terms of detection probability under a certain false alarm probability and sensing update rate in terms of the number of sensing measurement updates per time unit. It is numerically shown in [13] and [34], that maximizing the sensing SINR improves the target detection probability under a constant false alarm probability. When studying other sensing tasks, it is naturally desired to keep the sensing SINR above a required threshold denoted by $\gamma_{\rm s}$, i.e., $\mathsf{SINR}_{s}\geq\gamma_{\rm s}$. This fact motivates the optimization problems presented in Section VI. In the following subsections, we present sensing SINR and sensing update rate.
IV-A Sensing SINR
From [34] , the sensing SINR is given by
$\mathsf{SINR_{\rm s}}=\frac{\bm{\rho}^{T}\textbf{A}\bm{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}\bm{\rho}},$ (35)
where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$ and
[…]
×$\hskip 8.53581pt\times\textbf{W}_{k}\Bigg{)}\textbf{D}_{\rm s}[m],$ (36)
B $=\sum_{m=1}^{L_{d}}\textbf{D}_{\rm s}^{H}[m]\left(\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\right)$
×$\hskip 8.53581pt\times\textbf{D}_{\rm s}[m].$ (37)
@@ IV-A Sensing SINR
$\overline{\mathsf{SINR}}_{s}$ $=\frac{M\bm{\rho}^{T}\textbf{A}_{D}\bm{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}},$ (38)
@@
IV-B Sensing Update Rate (SUR)
In our system model, the sensing information is updated once throughout each transmission block. Thus, the sensing update rate, denoted as $R_{\rm s}$, is
@@ V E2E Power Consumption Modeling
$P_{\rm total}$ $=\underbrace{\sum_{k=1}^{N_{\rm tx}}\left(P_{\textrm{AP},0}^{\rm tx}+\Delta^{\rm tr}P_{k}\right)+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}}_{\textrm{radio site power consumption}}+P_{\rm cloud}$ (42)
where $\Delta^{\rm tr}$ is the slope of load-dependent transmission power consumption of each AP. The $P_{k}$ is the transmission power of the ISAC AP $k$ from (7). The power consumption at the cloud is modeled as
$P_{\rm cloud}$ $=P_{\rm fixed}+\frac{1}{\sigma_{\rm cool}}\Bigg{(}P_{\rm cloud,0}^{\rm proc}+\Delta_{\rm cloud}^{\rm proc}\frac{C_{\rm cloud}}{C_{\rm max}}\Bigg{)},$ (43)
where $P_{\rm fixed}$ is the fixed power consumption at the cloud which is independent of the load. $P_{\rm cloud,0}^{\rm proc}$ is the processing power consumption in the idle mode. $\sigma_{\rm cool}\in(0,1]$ and $\Delta_{\rm cloud}^{\rm proc}$ denote the cooling efficiency of the cloud and the slope of the load-dependent power consumption for processing at the digital unit (DU) in the cloud, respectively. Moreover, $C_{\rm max}$ and $C_{\rm cloud}\in[0,C_{\rm max}]$ are the maximum processing capacity of the processing resources and the total processing resource utilization in GOPS, respectively[30, 39]. The processing resource utilization can be expressed as
@@
In the following parts, the required GOPS for communication and sensing in our system model is computed, respectively. We assess the computational complexity where only the numbers of real multiplications and divisions are counted. Each complex multiplication is equal to four real multiplications. We also consider memory overhead in arithmetic operation calculations by multiplying each operation by two as done in [40, 30]. Hence, each complex multiplication is counted as $4\cdot 2=8$ operations in computing the total GOPS.
@@ V-A GOPS Analysis of Digital Operations for Communication at the Cloud
In this section, we analyze the GOPS for digital signal processing corresponding to the communication tasks including the uplink channel estimation and downlink transmission. To compute the number of real multiplications, we mainly follow the GOPS analysis in [36, App. B], [17].
Let $C_{\textrm{ch-est}}$ denote the computational complexity of the LMMSE channel estimation approach for all the APs. To compute the channel estimates, we first obtain the vectors $\textbf{y}_{t_{i},k}^{p}$. From [36, App. B], the multiplication of one matrix of size $M\times L_{p}$ with a vector of size $L_{p}\times 1$ results in $ML_{p}$ complex multiplications. Hence, obtaining $\textbf{y}_{t_{i},k}^{p}$ for $N_{\rm ue}$ UEs at all $N_{\rm tx}$ APs costs $8ML_{p}N_{\rm ue}N_{\rm tx}$ real multiplications/division in total, if $L_{p}\geq N_{\rm ue}$. Otherwise, the number of real multiplications/divisions would be $8ML_{p}^{2}N_{\rm tx}$. Moreover, we need to compute the matrices $\textbf{R}_{i,k}^{\prime}$ and $\bm{\Psi}_{t_{i},k}^{\prime}$, given that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ are known. However, this pre-computation can be neglected since the channel statistics are usually constant for a while and there is no need to compute them every coherence block. Finally, the multiplications in (10) costs $8M^{2}N_{\rm ue}N_{\rm tx}$ real operations. Then, $C_{\textrm{ch-est}}$ is equal to
@@
The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [36, App. B] is
@@
$=\left(12(MN_{\rm tx})^{2}+16MN_{\rm tx}\right)N_{\rm ue}$
$\quad+\frac{8\left((MN_{\rm tx})^{3}-MN_{\rm tx}\right)}{3}.$ (46)
Reciprocity calibration and multiplication of the symbols by the precoding vectors, each costs $8L_{d}MN_{\rm ue}N_{\rm tx}$ real operations [41, 30]. Multiplying by the power coefficients also costs $4L_{d}MN_{\rm ue}N_{\rm tx}$. Finally, the GOPS corresponding to communication processing (i.e., channel estimation, precoding and reciprocity calibration) is computed as
$C_{\textrm{proc}}^{\rm c}=\frac{B}{L10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+20L_{d}MN_{\rm ue}N_{\rm tx}\Bigg{)},$ (47)
where we divided the total giga operations by the coherence time $L/B$.

### v3

@@ II-B Multi-Static Sensing
We employ multi-static sensing, utilizing multiple transmit and receive APs within the network. Target sensing is performed during the downlink phase. A LOS connection is assumed to exist between the target location and each participating transmit and receive AP. In the presence of the target, each receive AP captures both the reflected signals from the target and undesired signals, referred to as clutter. The clutter, being independent of the target’s presence, is treated as interference for sensing purposes. Without loss of generality, we assume that the LOS paths between transmit and receive APs are known and can be effectively canceled out. Consequently, the interference signals correspond to the reflected paths through obstacles and are henceforth denoted as target-free channels.
@@
Let $\textbf{H}_{r,k}\in\mathbb{C}^{M\times M}$ denote the target-free channel matrix between transmit AP $k$ and receive AP $r$, which cause interference for sensing. We use the correlated Rayleigh fading model for the NLOS channels $\textbf{H}_{r,k}$, which is written as $\textbf{H}_{r,k}=\textbf{R}^{\frac{1}{2}}_{{\rm rx},(r,k)}\textbf{W}_{{\rm ch},{(r,k)}}\left(\textbf{R}^{\frac{1}{2}}_{{\rm tx},(r,k)}\right)^{T}$ using the Kronecker model [39]. The matrix $\textbf{W}_{{\rm ch},(r,k)}\in\mathbb{C}^{M\times M}$ is a random matrix with independent and identically distributed (i.i.d.) entries with $\mathcal{CN}(0,1)$ distribution. The matrix $\textbf{R}_{{\rm rx},(r,k)}\in\mathbb{C}^{M\times M}$ represents the spatial correlation matrix corresponding to receive AP $r$ and with respect to the direction of transmit AP $k$. Similarly, $\textbf{R}_{{\rm tx},(r,k)}\in\mathbb{C}^{M\times M}$ is the spatial correlation matrix corresponding to transmit AP $k$ and with respect to the direction of receive AP $r$. The channel gain is determined by the geometric path loss and shadowing, and is included in the spatial correlation matrices. The received signal at AP $r$ in the presence of the target and for $m\in\{1,\ldots,L_{d}\}$, can be formulated as
@@
$+\sum_{k=1}^{N_{\rm tx}}\underbrace{\textbf{H}_{r,k}\textbf{x}_{k}[m]}_{\textrm{clutter}}+\textbf{n}_{r}[m],$ (7)
where $\textbf{n}_{r}[m]\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}\textbf{I}_{M})$ is the receiver noise at the $M$ antennas of receive AP $r$. The second term in (II-B) refers to clutter which acts as interference for the target detection. Here, $\beta_{r,k}$ is the channel gain, including the path loss from transmit AP $k$ to receive AP $r$ through the target and the variance of bi-static radar cross-section (RCS) of the target denoted by $\sigma_{\rm rcs}$. $\beta_{r,k}$ is computed according to the radar range equation for bi-static sensing in [7, Chap. 2], given by
$\beta_{r,k}=\frac{\lambda_{c}^{2}\,\sigma_{\rm rcs}}{(4\pi)^{3}d_{{\rm tx},k}^{2}d_{{\rm rx},r}^{2}}.$ (8)
@@
Following the same notation as in [40], in (II-B), the known part of each reflected path is denoted by ${\textbf{g}_{r,k}[m]\in\mathbb{C}^{M}}$, defined as
@@
Each receive AP sends their respective signals $\textbf{y}_{r}[m]$, for $r=1,\ldots,N_{\rm rx}$, to the edge cloud to form the collective received signal $\textbf{y}[m]=\begin{bmatrix}\textbf{y}_{1}^{T}[m]&\ldots&\textbf{y}_{N_{\rm rx}}^{T}[m]\end{bmatrix}^{T}$, which can be expressed as
@@
$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\bm{\mathfrak{h}}+\textbf{n}[m]$
$=\textbf{G}[m]\bm{\alpha}+\textbf{X}[m]\bm{\mathfrak{h}}+\textbf{n}[m],$ (10)
where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\ldots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\ldots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\bm{\mathfrak{h}}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channel[40, Sec. V]. In addition, we have defined $\textbf{G}[m]=\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)$ and $\textbf{X}[m]=\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)$.
We employ two maximum a posteriori ratio test (MAPRT) detectors with two levels of complexity: the clutter-unaware detector and the clutter-aware detector, proposed in [12] and [40, Lem. 2], respectively. The clutter-unaware detector ignores the presence of clutter for the sake of complexity reduction, whereas the clutter-aware detector accounts for unknown clutter. The corresponding test statistics are given by
@@ III Reliability and Delay Analysis for URLLC
For most URLLC applications, short codewords are usually needed to satisfy latency constraints where codes with short blocklengths, e.g., 50-400 symbols are employed. Short blocklengths result in decoding errors; therefore, to ensure reliability, it is essential to consider the finite blocklength regime for modeling the decoding error probability (DEP) [41, 42, 8].
We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements. In this section, we derive an upper bound on the DEP and the transmission delay, both crucial aspects considered as URLLC requirements.
In the finite blocklength regime, the communication data cannot be transmitted without error. From [9], ergodic data rate of UE $i$ can be approximated as
$R_{i}\!\approx\!\mathbb{E}\Bigg{\{}\!\!\left(1\!-\!\beta\right)\!\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\!\sqrt{\!\frac{(1\!-\!\beta)V_{i}}{L}}\Bigg{\}},\vskip 0.0pt$ (17)
where $\beta=\frac{L_{p}}{L}$, $\epsilon_{i}$ denotes the DEP when transmitting $b_{i}$ bits to UE $i$, $\mathsf{SINR}^{\rm(dl)}_{i}$ is the instantaneous downlink communication SINR for UE $i$, $V_{i}=1-\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)^{-2}$ is the channel dispersion, and $Q(\cdot)$ refers to the Gaussian Q-function. Due to the fact that $V_{i}<1$, the ergodic data rate can be lower bounded by
@@
Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [34, Thm. 6.1] and [40, Lem. 1],
@@
where $D_{i}^{\rm(th)}$ is the maximum tolerable delay by UE $i$ and $D_{i}^{\rm(ub)}\leq D_{i}^{\rm(th)}$ should be satisfied to guarantee the delay requirement. This implies that the blocklength cannot exceed $D^{\rm(th)}_{i}B(1-\epsilon_{i}^{\rm(th)})$. Thus, we can define the maximum tolerable blocklength by $L_{\rm max,c}$, where
$L_{\rm max,c}=\min\left\{D^{\rm(th)}_{i}B\left(1-\epsilon_{i}^{\rm(th)}\right)|i=1,\ldots,N_{\rm ue}\right\}.$ (26)
@@ IV Sensing Analysis
We assess the sensing performance in terms of detection probability under a certain false alarm probability and refreshing rate in terms of the number of sensing measurement updates per time unit. Detection probability refers to the likelihood of correctly identifying the presence of a target, whereas false alarm probability denotes the likelihood of incorrectly detecting a target when none is present.
Maximizing the sensing SINR enhances the probability of target detection under a fixed false alarm probability [7, Chapters 3 and 15]. For other sensing tasks—such as estimating range, velocity, and similar parameters— it is also desired to maintain the sensing SINR above a required threshold denoted by $\gamma_{\rm s}$, i.e., $\mathsf{SINR}_{\rm s}\geq\gamma_{\rm s}$.This requirement serves as a key motivation for the optimization problem formulated in Section  VI. In this work, we consider a minimum sensing SINR and refreshing rate as the sensing requirements.
@@
$\overline{\mathsf{SINR}}_{\rm s}$ $=\frac{M\bm{\rho}^{T}\textbf{A}_{D}\bm{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}},$ (27)
@@
According to the 3GPP specification [6], refreshing rate is defined as the rate at which the sensing result is generated by the sensing system. It is the inverse of the time elapsed between two successive sensing results. In our system model, the sensing information is updated once throughout each transmission block. Thus, the sensing service latency is $\frac{L}{B}$ and the refreshing rate, denoted as $R_{\rm s}$, is
@@ V E2E Power Consumption Modeling
$P_{\rm total}$ $=\underbrace{\Delta^{\rm tr}P_{\rm tr}+\sum_{k=1}^{N_{\rm tx}}P_{\textrm{AP},0}^{\rm tx}+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}}_{\textrm{radio site power consumption}}+P_{\rm cloud}$ (31)
where $\Delta^{\rm tr}$ is the slope of load-dependent transmission power consumption of each AP. The average power consumption for transmission is given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\bm{\rho}\|^{2}$, where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$ due to the unit-power centralized precoding vectors. The power consumption at the cloud is modeled as
$P_{\rm cloud}$ $=P_{\rm fixed}+\frac{1}{\sigma_{\rm cool}}\Bigg{(}N_{\rm GPP}P_{\rm cloud,0}^{\rm proc}+\Delta_{\rm cloud}^{\rm proc}\frac{C_{\rm cloud}}{C_{\rm max}}\Bigg{)},$ (32)
where $P_{\rm fixed}$ is the fixed power consumption at the cloud, which is independent of the load. $N_{\rm GPP}=\lceil\frac{C_{\rm cloud}}{C_{\rm max}}\rceil$ is the number of required general-purpose processors (GPPs), and $P_{\rm cloud,0}^{\rm proc}$ is the processing power consumption in the idle mode. $\sigma_{\rm cool}\in(0,1]$ and $\Delta_{\rm cloud}^{\rm proc}$ denote the cooling efficiency of the cloud and the slope of the load-dependent power consumption for processing at the digital unit (DU) in the cloud, respectively. Moreover, $C_{\rm max}$ and $C_{\rm cloud}$ are the maximum processing capacity of the processing resources in each GPP and the total processing resource utilization in GOPS, respectively [18, 43]. The processing resource utilization can be expressed as
@@
In the following parts, the required GOPS for communication and sensing in our system model is computed, respectively. We assess the computational complexity where only the numbers of real multiplications and divisions are counted. Each complex multiplication is equal to four real multiplications. We also consider memory overhead in arithmetic operation calculations by multiplying each operation by two as done in [44, 18]. Hence, each complex multiplication is counted as 42=84$\cdot$ 2=84 ⋅ 2 = 8 operations in computing the total GOPS.
@@ V-A GOPS Analysis of Digital Operations for Communication at the Cloud
In this section, we analyze the GOPS for digital signal processing corresponding to the communication tasks including the uplink channel estimation and downlink transmission. To compute the number of real multiplications, we mainly follow the GOPS analysis in [37, App. B], [17].
Let $C_{\textrm{ch-est}}$ denote the computational complexity of the LMMSE channel estimation approach for all the APs. To compute the channel estimates, we first obtain the vectors $\textbf{y}_{t_{i},k}^{p}$. From [37, App. B], the multiplication of one matrix of size $M\times L_{p}$ with a vector of size $L_{p}\times 1$ results in $ML_{p}$ complex multiplications. Hence, obtaining $\textbf{y}_{t_{i},k}^{p}$ for $N_{\rm ue}$ UEs at all $N_{\rm tx}$ APs costs $8ML_{p}N_{\rm ue}N_{\rm tx}$ real multiplications/division in total, if $L_{p}\geq N_{\rm ue}$. Otherwise, the number of real multiplications/divisions would be $8ML_{p}^{2}N_{\rm tx}$. Moreover, we need to compute the matrices $\textbf{R}_{i,k}^{\prime}$ and $\bm{\Psi}_{t_{i},k}^{\prime}$, given that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ are known. However, this pre-computation can be neglected since the channel statistics are usually constant for a while and there is no need to compute them every coherence block. We use the LMMSE channel estimation approach in [38] which costs $8M^{2}N_{\rm ue}N_{\rm tx}$ real operations. Then, $C_{\textrm{ch-est}}$ is equal to
@@
The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [37, App. B] is
@@
$=\left(12(MN_{\rm tx})^{2}+16MN_{\rm tx}\right)N_{\rm ue}+\frac{8\left((MN_{\rm tx})^{3}-MN_{\rm tx}\right)}{3}.$ (35)
Reciprocity calibration and multiplication of the symbols by the precoding vectors, each costs $8L_{d}MN_{\rm ue}N_{\rm tx}$ real operations [45, 18]. Multiplying by the power coefficients also costs $4L_{d}MN_{\rm ue}N_{\rm tx}$. Finally, the GOPS corresponding to communication processing (i.e., channel estimation, precoding and reciprocity calibration) is computed as in (36) where we divided the total giga operations by the coherence time $L/B$.
$C_{\textrm{proc}}^{\rm c}=\frac{B}{L10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+20L_{d}MN_{\rm ue}N_{\rm tx}\Bigg{)}.$ (36)
TABLE II: COMPUTATIONAL COMPLEXITY FOR CLUTTER-AWARE DETECTOR
Operation Computational complexity
$\textbf{G}[m],\quad\forall m$ $20\,L_{d}\,MN_{\rm tx}N_{\rm rx}$
a $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,$
b $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}$
C $4\,L_{d}MN_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)$
D $4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)$
E $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}^{2}$
$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$
$\frac{8}{3}\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}\right)$
$\textbf{D}^{-1}$ $\frac{8}{3}\left((M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}\right)$
$\begin{bmatrix}\textbf{a}^{H}\!&\textbf{b}^{H}\end{bmatrix}\!\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$
$8\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+8(1+M^{2})N_{\rm tx}N_{\rm rx}$